# From multitype branching Brownian motions to branching Markov additive processes

### v1

@@ 5.1 The spine decomposition with respect to the additive martingale
$\mathrm{d}\widetilde{\mathbf{P}}_{x,i}(\tau,M,\xi)\big|_{\widetilde{\mathcal{F}}_{t}}$ (5.1)
$=$ $\mathrm{d}\mathbb{P}_{x,i}\left((X_{\xi},J_{\xi})_{t}\right)\mathrm{d}L^{\beta(J_{\xi})}(\mathbf{n}_{t})\prod_{v\prec\xi_{t}}\mu_{J_{\xi}(d_{v})}(A_{v})\prod_{v\prec\xi_{t}}\left[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{X_{\xi}(d_{v}),J_{\xi}(d_{v})}((\tau,M)_{t-d_{v}}^{v,j})\right],$ (5.2)
@@
To prove that $(\zeta_{t},t\geq 0,\widetilde{\mathbf{P}}_{x,i})$ is a martingale, we proceeds by decomposing $\zeta_{t}$ into the product of three parts $\zeta_{t}^{(1)}$, $\zeta_{t}^{(2)}$ and $\zeta_{t}^{(3)}$, which will be defined sequentially as the argument develops. We also need the following definition from [Liu11].
@@
First, by [Hardy09, Theorem 5.4], we have the following lemma.
@@
Recall that $W_{\theta}(t)=\sum_{u\in\mathcal{N}_{t}}e^{-\left(\theta X_{u}(t)+\lambda(\theta)t\right)}V_{J_{u}(t)}(\theta)$. By the many-to-one formula (for example, see [Harris-Roberts17]), we have
@@
The next lemma follows from [Hardy09, Theorem 5.5].
@@
$\widetilde{\mathbf{P}}_{x,i}(\zeta_{t}\mid\mathcal{F}_{t})=\sum_{u\in\mathcal{N}_{t}}\prod_{v\prec u}A_{v}e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta)\widetilde{\mathbf{P}}_{x,i}(\mathbbm{1}_{\left\{\xi_{t}=u\right\}}\mid\mathcal{F}_{t})=\sum_{u\in\mathcal{N}_{t}}e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta),$ (5.14)
@@
$\mathrm{d}\widetilde{\mathbf{P}}^{\theta}_{x,i}(\tau,M,\xi)|_{\widetilde{\mathcal{F}}_{t}}$ (5.16)
$=$ $\frac{\zeta_{t}^{(2)}}{\zeta_{0}}\mathrm{d}\mathbb{P}_{x,i}\left((X_{\xi},J_{\xi})_{t}\right)\zeta_{t}^{(1)}\mathrm{d}L^{\beta(J_{\xi})}(\mathbf{n}_{t})\zeta_{t}^{(3)}\prod_{v\prec\xi_{t}}\mu_{J_{\xi}(d_{v})}(A_{v})\prod_{v\prec\xi_{t}}\left[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{X_{\xi}(d_{v}),J_{\xi}(d_{v})}((\tau,M)_{t-d_{v}}^{v,j})\right]$ (5.17)
$=$ $\mathrm{d}\mathbb{P}_{x,i}^{\theta}\left((X_{\xi},J_{\xi})_{t}\right)\mathrm{d}L^{(m\beta)(J_{\xi})}(\mathbf{n}_{t})\prod_{v\prec\xi_{t}}\frac{\mu_{J_{\xi}(d_{v})}(A_{v})A_{v}}{m(J_{\xi}(d_{v}))}\prod_{v\prec\xi_{t}}\left[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{X_{\xi}(d_{v}),J_{\xi}(d_{v})}((\tau,M)_{t-d_{v}}^{v,j})\right].$ (5.18)
@@
Now we prove Theorem 2.7, allowing the possibility of no offspring when a particle dies. Our proof follows the construction of the spine decomposition for branching Markov processes given in [RS20].
We now require a slight modification to the definition of a marked tree with a distinguished spine. Let $\dagger$ be a fictitious node not in $\tau$. Following the construction in [RS20, Page 6], a spine $\xi$ on a marked tree $(\tau,M)$ is a subset of $\tau\cup\{\dagger\}$ such that
@@
Then we call $d_{\dagger-1}$ the “lifetime” of the spine. Let $\xi_{t}:=u$ be the unique element $u\in\xi\cap(N_{t}\cup\{\dagger\})$. Similar to [RS20, Equation (2.2)], we have
$\mathrm{d}\widetilde{\mathbf{P}}_{x,i}(\tau,M,\xi)\big|_{\widetilde{\mathcal{F}}_{t}}$ (5.19)
$=$ $\mathbbm{1}_{\left\{\xi_{t}\in\tau\right\}}\mathrm{d}\mathbb{P}_{x,i}\left((X_{\xi},J_{\xi})_{t}\right)\mathrm{d}L^{\beta(J_{\xi})}(\mathbf{n}_{t})\prod_{v\prec\xi_{t}}\mu_{J_{\xi}(d_{v})}(A_{v})\prod_{v\prec\xi_{t}}\left[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{X_{\xi}(d_{v}),J_{\xi}(d_{v})}((\tau,M)_{t-d_{v}}^{v,j})\right]$ (5.20)
$+\mathbbm{1}_{\left\{\xi_{t}=\dagger\right\}}\mathrm{d}\mathbb{P}_{x,i}\left((X_{\xi},J_{\xi})_{t}\right)\mathrm{d}L^{\beta(J_{\xi})}(\mathbf{n}_{t})\prod_{v\prec\dagger-1}\mu_{J_{\xi}(d_{v})}(A_{v})\prod_{v\prec\dagger-1}\left[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{X_{\xi}(d_{v}),J_{\xi}(d_{v})}((\tau,M)_{t-d_{v}}^{v,j})\right].$ (5.21)
@@
$\zeta_{t}^{(1)}:=\prod_{v<\xi_{t}}m(J_{\xi}(d_{v}))\cdot\exp\left\{-\int_{0}^{t\wedge d_{\dagger-1}}((m-1)\beta)(J_{\xi}(s))\mathrm{d}s\right\},$ (5.23)
@@
$\zeta_{t}^{(3)}:=\prod_{v\prec\xi_{t}}\frac{A_{v}}{m(J_{\xi}(d_{v}))}=\mathbbm{1}_{\left\{\xi_{t}\in\mathcal{N}_{t}\right\}}\prod_{v\prec\xi_{t}}\frac{A_{v}}{m(J_{\xi}(d_{v}))},$ (5.25)
where the last equality follows from the fact that $A_{\dagger-1}=0$ when $\xi_{t}=\dagger$. By [RS20, Lemma 2.6], we know that Lemmas 5.2 and 5.4 hold in this case. By Lemma 2.6, it is easy to show that $\{\zeta_{t}^{(2)},t\geq 0\}$ is a $\mathbf{P}_{x,i}$-martingale with respect to $\widetilde{\mathcal{F}}_{t}$. Then, we have
$\zeta_{t}=\zeta_{t}^{(1)}\zeta_{t}^{(2)}\zeta_{t}^{(3)}$ (5.26)
and Lemma 2.5 holds. The definition of $\zeta_{t}$ yields that $\widetilde{\mathbf{P}}_{x,i}^{\theta}(\xi_{t}\in\mathcal{N}_{t})=1$. Thus,
$\,\mathrm{d}\widetilde{\mathbf{P}}^{\theta}_{x,i}(\tau,M,\xi)|_{\widetilde{\mathcal{F}}_{t}}$ (5.27)
$=$ $\,\mathrm{d}\mathbb{P}_{x,i}^{\theta}\left((X_{\xi},J_{\xi})_{t}\right)\mathrm{d}L^{(m\beta)(J_{\xi})}(\mathbf{n}_{t})\prod_{v\prec\xi_{t}}\frac{\mu_{J_{\xi}(d_{v})}(A_{v})A_{v}}{m(J_{\xi}(d_{v}))}\prod_{v\prec\xi_{t}}\left[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{(X_{\xi}(d_{v}),J_{\xi}(d_{v}))}((\tau,M)_{t-d_{v}}^{v,j})\right].$ (5.28)
@@
We use the method of extended generator in [PalRol2002]. Following [PalRol2002] and the notations in Proposition 1.2, we first decompose $\chi_{t}=\chi_{t}^{(1)}+\chi^{(2)}_{t}$, where $\{\chi_{t}^{(1)},t\geq 0\}$ and $\{\chi^{(2)}_{t},t\geq 0\}$ are two independent processes,
@@
$\mathbf{A}^{(1)}f(x,i)=\sum_{k\neq i}q_{ik}\int_{-\infty}^{\infty}(f(x+y,k)-f(x,i))\mathbb{P}(U_{ik}\in\mathrm{d}y)=\sum_{k\in\mathcal{I}}q_{ik}\int_{-\infty}^{\infty}f(x+y,k)\mathbb{P}(U_{ik}\in\mathrm{d}y),$
@@
By [PalRol2002, Lemma 3.1], $(E^{\widetilde{h}}(t),t\geq 0)$ is a $\mathbb{P}_{x,i}$-local martingale. Since $\mathbb{E}_{x,i}[E^{\widetilde{h}}(t)]=\mathbb{E}_{x,i}[\frac{\Xi_{\theta}(t)}{\Xi_{\theta}(0)}]=1$ by (5.9), then $E^{\widetilde{h}}$ is a true martingale. According to [PalRol2002, Lemma 4.1][PalRol2002, Theorem 4.2], define the probability change
@@
$\widetilde{\mathbf{A}}^{(1)}f(x,i)$ $:=\sum_{k\neq i}\widetilde{q}_{ik}\int_{-\infty}^{\infty}(f(x+y,k)-f(x,i))\mathbb{P}(\widetilde{U}_{ik}\in\mathrm{d}y)=\sum_{k\in\mathcal{I}}\widetilde{q}_{ik}\int_{-\infty}^{\infty}f(x+y,k)\mathbb{P}(\widetilde{U}_{ik}\in\mathrm{d}y).$
@@
$\widetilde{\mathbf{A}}F(x,y,i)$ $=g(x)(\widetilde{\mathbf{A}}^{(1)}f)(y,i)+f(y,i)(\widetilde{\mathbf{A}}^{i}g)(x)+f(y,i)g(x)((q_{ii}-\widetilde{q}_{ii})+\phi_{i}(\theta)-\lambda(\theta)+\beta_{i}(m_{i}-1))$ (5.40)
@@
This implies that, under $\mathbb{P}_{x,i}^{\theta}$, $(\chi,\Theta)$ is a MAP with the given characteristics in the lemma. ∎

### v2

@@ 5.1 The spine decomposition with respect to the additive martingale
$\mathrm{d}\widetilde{\mathbf{P}}_{x,i}(\tau,M,\xi)\big|_{\widetilde{\mathcal{F}}_{t}}$ $=\mathrm{d}\mathbb{P}_{x,i}\left((X_{\xi},J_{\xi})_{t}\right)\mathrm{d}L^{\beta(J_{\xi})}(\mathbf{n}_{t})\prod_{v\prec\xi_{t}}\mu_{J_{\xi}(d_{v})}(A_{v})$ (5.1)
$\qquad\times\prod_{v\prec\xi_{t}}\bigg[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{X_{\xi}(d_{v}),J_{\xi}(d_{v})}((\tau,M)_{t-d_{v}}^{v,j})\bigg],$ (5.2)
@@
To prove that $(\zeta_{t},t\geq 0,\widetilde{\mathbf{P}}_{x,i})$ is a martingale, we proceeds by decomposing $\zeta_{t}$ into the product of three parts $\zeta_{t}^{(1)}$, $\zeta_{t}^{(2)}$ and $\zeta_{t}^{(3)}$, which will be defined sequentially as the argument develops. We also need the following definition from [40].
@@
First, by [22, Theorem 5.4], we have the following lemma.
@@
Recall that $W_{\theta}(t)=\sum_{u\in\mathcal{N}_{t}}e^{-\left(\theta X_{u}(t)+\lambda(\theta)t\right)}V_{J_{u}(t)}(\theta)$. By the many-to-one formula (for example, see [23]), we have
@@
The next lemma follows from [22, Theorem 5.5].
@@
$\widetilde{\mathbf{P}}_{x,i}(\zeta_{t}\mid\mathcal{F}_{t})$ $=\sum_{u\in\mathcal{N}_{t}}\prod_{v\prec u}A_{v}e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta)\widetilde{\mathbf{P}}_{x,i}(\mathbbm{1}_{\left\{\xi_{t}=u\right\}}\mid\mathcal{F}_{t})$ (5.14)
$=\sum_{u\in\mathcal{N}_{t}}e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta),$ (5.15)
@@
$\mathrm{d}\widetilde{\mathbf{P}}^{\theta}_{x,i}(\tau,M,\xi)|_{\widetilde{\mathcal{F}}_{t}}=$ $\,\frac{\zeta_{t}^{(2)}}{\zeta_{0}}\mathrm{d}\mathbb{P}_{x,i}\left((X_{\xi},J_{\xi})_{t}\right)\zeta_{t}^{(1)}\mathrm{d}L^{\beta(J_{\xi})}(\mathbf{n}_{t})\zeta_{t}^{(3)}\prod_{v\prec\xi_{t}}\mu_{J_{\xi}(d_{v})}(A_{v})$ (5.17)
$\,\prod_{v\prec\xi_{t}}\Bigg[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{X_{\xi}(d_{v}),J_{\xi}(d_{v})}((\tau,M)_{t-d_{v}}^{v,j})\Bigg]$ (5.18)
$=$ $\,\mathrm{d}\mathbb{P}_{x,i}^{\theta}\left((X_{\xi},J_{\xi})_{t}\right)\mathrm{d}L^{(m\beta)(J_{\xi})}(\mathbf{n}_{t})\prod_{v\prec\xi_{t}}\frac{\mu_{J_{\xi}(d_{v})}(A_{v})A_{v}}{m(J_{\xi}(d_{v}))}$ (5.19)
$\,\prod_{v\prec\xi_{t}}\Bigg[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{X_{\xi}(d_{v}),J_{\xi}(d_{v})}((\tau,M)_{t-d_{v}}^{v,j})\Bigg].$ (5.20)
@@
Now we prove Theorem 2.7, allowing the possibility of no offspring when a particle dies. Our proof follows the construction of the spine decomposition for branching Markov processes given in [50].
We now require a slight modification to the definition of a marked tree with a distinguished spine. Let $\dagger$ be a fictitious node not in $\tau$. Following the construction in [50, Page 6], a spine $\xi$ on a marked tree $(\tau,M)$ is a subset of $\tau\cup\{\dagger\}$ such that
@@
Then we call $d_{\dagger-1}$ the “lifetime” of the spine. Let $\xi_{t}:=u$ be the unique element $u\in\xi\cap(N_{t}\cup\{\dagger\})$. Define $D_{t}:=\{u\in\tau:d_{u}\leq t,A_{u}=0\}$ be the set of particles that died, before or at time $t$, with no offspring. For a particle $u\in D_{t}$, two distinct cases arise, each requiring a separate treatment: either $m(J_{u}(d_{u}))=0$, or $m(J_{u}(d_{u}))>0$ but $A_{u}=0$. Moreover, if $\xi_{t}=\dagger$, then there is a unique $u\in D_{t}$ such that $\dagger-1=u$. Similar to [50, Equation (2.2)], we have
$\mathrm{d}$ $\widetilde{\mathbf{P}}_{x,i}(\tau,M,\xi)\big|_{\widetilde{\mathcal{F}}_{t}}=\mathbbm{1}_{\left\{\xi_{t}\in\tau\right\}}\mathrm{d}\mathbb{P}_{x,i}\left((X_{\xi},J_{\xi})_{t}\right)\mathrm{d}L^{\beta(J_{\xi})}(\mathbf{n}_{t})\prod_{v\prec\xi_{t}}\mu_{J_{\xi}(d_{v})}(A_{v})$ (5.21)
$\quad\times\prod_{v\prec\xi_{t}}\Bigg[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{X_{\xi}(d_{v}),J_{\xi}(d_{v})}((\tau,M)_{t-d_{v}}^{v,j})\Bigg]+\mathbbm{1}_{\left\{\xi_{t}=\dagger\right\}}\mathrm{d}\mathbb{P}_{x,i}\left((X_{\xi},J_{\xi})_{t}\right)\mathrm{d}L^{\beta(J_{\xi})}(\mathbf{n}_{t})$ (5.22)
$\quad\times\prod_{v\prec\dagger-1}\mu_{J_{\xi}(d_{v})}(A_{v})\prod_{v\prec\dagger-1}\Bigg[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{X_{\xi}(d_{v}),J_{\xi}(d_{v})}((\tau,M)_{t-d_{v}}^{v,j})\Bigg].$ (5.23)
@@
$\zeta_{t}:=\sum_{u\in\mathcal{N}_{t}}\bigg(\prod_{v\prec u}A_{v}\bigg)e^{-\theta X_{u}(t)-\lambda(\theta)t}V_{J_{u}(t)}(\theta)\mathbbm{1}_{\left\{\xi_{t}=u\right\}}.$ (5.24)
@@
$\zeta_{t}^{(1)}:$ $=\prod_{v\prec\xi_{t}}m(J_{\xi}(d_{v}))\cdot\exp\left\{-\int_{0}^{t\wedge d_{\dagger-1}}((m-1)\beta)(J_{\xi}(s))\mathrm{d}s\right\},$ (5.25)
@@
$\zeta_{t}^{(3)}:$ $=\prod_{v\prec\xi_{t}}\frac{A_{v}}{m(J_{\xi}(d_{v}))}.$ (5.27)
When $m(J_{\xi}(d_{v}))=0$, it holds that $A_{v}=0$ a.s. and we use the convention $\frac{0}{0}=1$ such that $\frac{A_{v}}{m(J_{\xi}(d_{v}))}\mathbbm{1}_{\left\{m(J_{\xi}(d_{v}))=0\right\}}=\mathbbm{1}_{\left\{m(J_{\xi}(d_{v}))=0\right\}}$ a.s. In other words, we have
$\zeta_{t}^{(3)}$ $=\prod_{v\prec\xi_{t}}\bigg(\frac{A_{v}}{m(J_{\xi}(d_{v}))}\mathbbm{1}_{\left\{m(J_{\xi}(d_{v}))\neq 0\right\}}+\mathbbm{1}_{\left\{m(J_{\xi}(d_{v}))=0\right\}}\bigg)$ (5.28)
$=\mathbbm{1}_{\left\{\xi_{t}\in\mathcal{N}_{t}\right\}}\prod_{v\prec\xi_{t}}\frac{A_{v}}{m(J_{\xi}(d_{v}))}+\mathbbm{1}_{\left\{\xi_{t}=\dagger\right\}}\mathbbm{1}_{\left\{m(J_{\xi}(d_{\dagger-1}))=0\right\}}\prod_{v\prec\dagger-1}\frac{A_{v}}{m(J_{\xi}(d_{v}))},$ (5.29)
where the last equality follows from the fact that $A_{\dagger-1}=0$ when $\xi_{t}=\dagger$ and $m(J_{\xi}(d_{\dagger-1}))>0$. Therefore, if $\xi_{t}=\dagger$, then either $m(J_{\xi}(d_{\dagger-1}))=0$ such that $\zeta^{(1)}_{t}=0$, or $m(J_{\xi}(d_{\dagger-1}))>0$ such that $\zeta^{(3)}_{t}=0$. This yields that $\zeta_{t}=0$ on $\{\xi_{t}=\dagger\}$. Then we have the identity
$\zeta_{t}=\zeta_{t}^{(1)}\zeta_{t}^{(2)}\zeta_{t}^{(3)},\qquad t\geq 0.$ (5.30)
From the discussion about $\zeta_{t}^{(3)}$ above we deduce Lemma 5.4 in this case. According to [12, Example 5.5.5], Lemma 5.2 still holds even when $m_{j}=0$ for some $j\in\mathcal{I}$. By Lemma 2.6, it is immediate that $\{\zeta_{t}^{(2)},t\geq 0\}$ is a $\mathbf{P}_{x,i}$-martingale with respect to $\widetilde{\mathcal{F}}_{t}$. We conclude that Lemma 2.5 holds. The definition of $\zeta_{t}$ yields that $\widetilde{\mathbf{P}}_{x,i}^{\theta}(\xi_{t}\in\mathcal{N}_{t})=1$ for any $t\geq 0$, which implies that $\widetilde{\mathbf{P}}_{x,i}^{\theta}(\xi_{t}\in\mathcal{N}_{t},\forall t\geq 0)=1$. Therefore,
$\mathrm{d}\widetilde{\mathbf{P}}^{\theta}_{x,i}(\tau,M,\xi)|_{\widetilde{\mathcal{F}}_{t}}=$ $\,\mathrm{d}\mathbb{P}_{x,i}^{\theta}\left((X_{\xi},J_{\xi})_{t}\right)\mathrm{d}L^{(m\beta)(J_{\xi})}(\mathbf{n}_{t})\prod_{v\prec\xi_{t}}\frac{\mu_{J_{\xi}(d_{v})}(A_{v})A_{v}}{m(J_{\xi}(d_{v}))}$ (5.31)
$\,\prod_{v\prec\xi_{t}}\Bigg[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{(X_{\xi}(d_{v}),J_{\xi}(d_{v}))}((\tau,M)_{t-d_{v}}^{v,j})\Bigg].$ (5.32)
@@
We use the method of extended generator in [47]. Following [47] and the notations in Proposition 1.2, we first decompose $\chi_{t}=\chi_{t}^{(1)}+\chi^{(2)}_{t}$, where $\{\chi_{t}^{(1)},t\geq 0\}$ and $\{\chi^{(2)}_{t},t\geq 0\}$ are two independent processes,
@@
$\mathbf{A}^{(1)}f(x,i)$ $=\sum_{k\neq i}q_{ik}\int_{-\infty}^{\infty}(f(x+y,k)-f(x,i))\mathbb{P}(U_{ik}\in\mathrm{d}y)$ (5.33)
$=\sum_{k\in\mathcal{I}}q_{ik}\int_{-\infty}^{\infty}f(x+y,k)\mathbb{P}(U_{ik}\in\mathrm{d}y),$ (5.34)
@@
By [47, Lemma 3.1], $(E^{\widetilde{h}}(t),t\geq 0)$ is a $\mathbb{P}_{x,i}$-local martingale. Since $\mathbb{E}_{x,i}[E^{\widetilde{h}}(t)]=\mathbb{E}_{x,i}[\frac{\Xi_{\theta}(t)}{\Xi_{\theta}(0)}]=1$ by (5.9), then $E^{\widetilde{h}}$ is a true martingale. According to [47, Lemma 4.1 and Theorem 4.2], define the probability change
@@
$\widetilde{\mathbf{A}}^{(1)}f(x,i)$ $:=\sum_{k\neq i}\widetilde{q}_{ik}\int_{-\infty}^{\infty}(f(x+y,k)-f(x,i))\mathbb{P}(\widetilde{U}_{ik}\in\mathrm{d}y)$
$=\sum_{k\in\mathcal{I}}\widetilde{q}_{ik}\int_{-\infty}^{\infty}f(x+y,k)\mathbb{P}(\widetilde{U}_{ik}\in\mathrm{d}y).$
@@
$\hskip 15.00002pt\widetilde{\mathbf{A}}F(x,y,i)$ (5.46)
$=g(x)(\widetilde{\mathbf{A}}^{(1)}f)(y,i)+f(y,i)(\widetilde{\mathbf{A}}^{i}g)(x)+f(y,i)g(x)((q_{ii}-\widetilde{q}_{ii})+\phi_{i}(\theta)-\lambda(\theta)+\beta_{i}(m_{i}-1))$ (5.47)
@@
This implies that, under $\mathbb{P}_{x,i}^{\theta}$, $(\chi,\Theta)$ is a MAP with characteristics given in the lemma. ∎